OpenStudy (anonymous):

Write an equation of a line in slope intercept form that is perpendicular to y = -4x -2 and passes through the point (-16, -11). Write an equation of a line in slope intercept form that is parallel to the line x + 4y = 6 and passes through (-8, 5). Write an equation of a line in slope intercept form that is perpendicular to the line 2x -3y = 12 and passes through the point (2, 6).

OpenStudy (nubeer):

ok few things u need to know.. when lines are parallel, they have the same slope.. and when one line is perpendicular to other line their slpe have this relation : m1*m2 = -1

OpenStudy (anonymous):

Alright! I know how to do them, but these are the last 3 questions to my study guide, I've been doing it for about two hours. If you could solve them I would be so grateful. Thanks! :)

OpenStudy (nubeer):

well first it's a learning site not a solution site.. everyone can walk you through the questions.. for 1st question.. can u tell me what will be the slope of the equation u have to find?

OpenStudy (amriju):

nubeer is correct....you need to determine the slope of the required line first...for that as u can understand u must bring the eqn of the given line to y=mx+c form....now notice if the required line is parellel or perp to the given line...then do as nubeer said and determine the slope...now slope =(y-Y)?(x-X)...where Y,X are the co-ords of the givn point through which the reqd line passes... q1: clearly slope of the given line=-4 slope of reqd line then...=1/4( note they r perp) line passes through (-16,-11) then eqn..: 1/4=(y-(-11))/(x-(-16)) thats x-4y=18 Q2: given line: x+4y=6 thats y=-x/4+6/4 this is the slope intercept form of the GIVEN line then slope of given line=-1/4 now slope of reqd line =-1/4( note they r parrellel) line passes thru(-8,5) eqn of reqd line: -1/4=(y-5)/(x-(-8)) thats x+4y=12 are these two ans correct...then u can do the thrd one...its the same as the first